find the equation of the line whose slope is 2 and sum of intercepts is a+b=6

Grade:11

1 Answers

Naren Ks
8 years ago
draw a simple figure appropriate to the question.
given :    x co-ordinate=(a,0) [y value on x axis is 0]      y co-ordinate=(0,b) [ x value on y axis is 0]   slope of the line =2
to find:  equation of the line
sol:   b-0\0-a=2 [slope formula when two points are given] = -b/a =2
solve the equations -b/a=2 and a+b=6 to find the value of a and b .substitute the co ordinates in the slope point form to get the equation of the required line